### v2

# Training PINNs with Hard Constraints and Adaptive Weights: An Ablation Study

Abstract
Training Physics-Informed Neural Networks (PINNs) to solve stiff time-dependent partial differential equations (PDEs) remains a significant challenge. The main difficulties lie in precisely enforcing initial conditions and balancing the various loss components. In stiff PDEs, where solutions show rapid transitions or sharp gradients, this balance becomes especially difficult. The wide range of scales in the dynamics can cause certain losses to dominate, leading to unstable or inefficient training. This paper presents a comprehensive ablation study focused on two pivotal training schemes: the enforcement of hard constraints for initial and boundary conditions, and the implementation of adaptive weights. We specifically examine their impact on the performance of PINNs when applied to stiff time-dependent PDEs from materials science and mathematical biology applications. We conduct extensive numerical experiments across a diverse range of time-dependent PDEs from Allen-Cahn, Cahn-Hillard, to Gray-Scott systems. We further discuss the implications of our findings for improving the robustness and efficiency of PINN training, particularly in settings where accurate representation of initial conditions and balanced loss contributions are paramount.
1 Introduction
Time-dependent partial differential equations (PDEs) constitute the mathematical foundation of numerous scientific and engineering disciplines, serving as essential tools for modeling a diverse range of physical phenomena. Their applications span from the intricate dynamics of fluid flow, as described by the Navier–Stokes equations, to the complex interplay of chemical reactions in biological systems, captured by reaction–diffusion models. Time-dependent PDEs offer a powerful framework for understanding and predicting the evolution of systems over time whitham2011linear ; burgers1948mathematical ; allen1979microscopic ; shen2010numerical ; bazant2017thermodynamic ; horstmann2013precipitation ; tian2015electrochemically ; hyman2014liquid ; takatori2015towards ; cahn1958free ; lee2014physical ; miranville2017cahn ; kudryashov1990exact ; michelson1986steady ; de1991turing ; nishiura1999skeleton ; zhabotinsky2007belousov ; cassani2021belousov ; kato1987nonlinear ; kevrekidis2001discrete ; fibich2015nonlinear . In particular, the imposition of periodic boundary conditions allows for the simulation of phenomena over effectively infinite physical domains, providing valuable insights into the intrinsic behavior of physical systems. However, even the numerical solution of these PDEs, especially in scenarios involving stiffness or complex non-linearities, can be exceedingly challenging, often demanding significant computational resources and specialized numerical techniques.
The advent of machine learning has ushered in a new era of scientific computing, offering powerful tools for data-driven modeling and analysis. Machine learning methodologies have shown remarkable success in various science/engineering fields, revolutionizing areas such as protein folding prediction PF2019 , drug discovery and development CARRACEDOREBOREDO20214538 , and even optimizing fundamental algorithmic processes such as matrix-vector multiplication MMML2022 . However, traditional machine learning approaches, which rely on vast datasets, can prove impractical in scientific contexts where data acquisition is expensive, time consuming, or simply infeasible. Furthermore, purely data-driven models often lack physical fidelity and model interpretability, as they may fail to capture the underlying physical principles governing the systems they aim to describe. Physics-informed machine learning (PIML), specifically physics-informed neural networks (PINNs), address these limitations by seamlessly integrating physical laws into the learning process, significantly reducing the amount of training data needed. By embedding constraints derived from the governing PDEs directly into the neural network architecture, PINNs enable accurate predictions even with sparse and noisy data yang2019physics ; chen2021solving ; raissi2018hidden ; raissi2019physics ; liu2019multi ; liu2021dual ; jin2021nsfnets ; cai2021deepm ; rad2020theory ; zhu2021machine ; chen2022using ; wang2023pinnsbased . Despite the promising applications, effective training of PINNs, especially for stiff time-dependent PDEs, remains a significant challenge. The inherent difficulty in propagating information from initial and boundary data points into the computational domain, coupled with the need to balance multiple loss terms, often leads to convergence issues and suboptimal solutions wang2020understanding ; wang2020and ; mcclenny2020self ; coutinho2022physics .
This paper presents an ablation study of training algorithms for PINNs applied to stiff partial differential equations (PDEs), focusing specifically on two schemes: hard constraints and self-adaptive loss weights mcclenny2020self . The hard-constraint scheme directly embeds information about initial and boundary conditions into the neural network architecture. This transformation reduces the effects of spectral bias from training various kinds of loss terms, leading to improved solution accuracy wang2020and . PINN’s training can treated as a regularized regression, where initial and boundary conditions, along with observational data, serve as training data, and the physics-informed loss acts as regularization. However, the multi-objective nature of the loss function and the sensitivity of PDE solutions to initial conditions can significantly hinder the search for global minimizers. By explicitly incorporating boundary and initial condition data into the network transformation, we streamline the training process, particularly for stiff time-dependent PDEs, where accurately capturing the initial conditions is critical. In addition, we investigate the use of adaptive loss weights to alleviate the spectral bias from the multi-objective loss in PINN training mcclenny2020self . Through our ablation study of these training schemes, we arrive at the conclusion that hard constrain transformation provides better improvement. We emphasize that hard constraints and self-adaptive weights are not the only effective strategies for handling stiff PDEs. These methods can be integrated with other advanced PINN techniques, such as causal PINNs wang2020and , time-marching PINNs wight2020solving , residual-based-attention (RBA) PINNs ANAGNOSTOPOULOS2024116805 , curriculum training krishnapriyan2021characterizing , and co-training PINNs zhong2024label , to achieve synergistic improvements in prediction accuracy and robustness. Through a rigorous ablation study, we demonstrate the efficacy of the proposed approach across a range of stiff time-dependent PDEs, offering valuable insights into optimal training strategies for PINNs in these challenging scenarios.
1.1 Related Work
PINNs have shown significant potential to solve a wide range of PDEs. However, challenges arise, particularly when dealing with “stiff” PDEs, where explicit numerical methods require extremely small time steps for stability. A primary concern is the imbalance between the fitting for initial data and residual components within the PINN loss function wight2020solving ; shin2020convergence ; wang2020understanding ; wang2020and . This imbalance often leads to a gradient descent that favors minimization of residual loss over data fitting loss, hindering convergence to accurate solutions. This issue is exacerbated in forward problems, where data is primarily concentrated at initial and boundary conditions, leaving limited information within the PDE domain. Specifically, in time-dependent problems, PINNs struggle to propagate information from initial conditions to subsequent time steps, a phenomenon widely documented in the literature wight2020solving ; mcclenny2020self ; krishnapriyan2021characterizing ; wang2022respecting ; haitsiukevich2022improved .
Several strategies have been proposed to address these challenges. One approach, introduced by Wight and Zhao wight2020solving , involves time-marching, where the time domain is divided into smaller intervals, and PINNs are sequentially trained on each interval, starting from the initial condition. While effective, this method can be computationally expensive due to the need to train multiple networks. To improve information propagation, researchers have emphasized the importance of weighting initial condition data and residual points near the initial time wight2020solving ; mcclenny2020self . McClenny and Konstantinidis mcclenny2020self demonstrated that their self-adaptive PINNs can autonomously learn these weights during training, particularly in scenarios with complex initial conditions, such as the wave equation.
Further advancements include the “sequence-to-sequence” approach proposed by Krishnapriyan et al. krishnapriyan2021characterizing , which sequentially learns each time step in a time-marching scheme. Wang et al. wang2022respecting identified the lack of “causality” in standard PINN training and proposed a corrective weighting scheme that gradually prioritizes data from early to later times, leveraging information from the PDE residue. This approach shares similarities with the self-adaptive weighting mechanism of McClenny and Konstantinidis mcclenny2020self . Additionally, Haitsiukevich et al. haitsiukevich2022improved introduced a pseudo-label approach, resembling self-training, to further enhance PINN convergence and accuracy. In zhong2024label , a mixed training scheme using labal-propagation between data points and collocation points is introduced and tested. For stiff parabolic PDEs such as viscous Burgers with small viscosity, the authors in coutinho2022physics proposed an data-driven artificial viscosity scheme to produce stable solutions and accurate to capture sharp transitions in the solutions.
Beyond loss weighting and time marching, structural modifications to PINNs have also been explored. Braga-Neto et al. braganeto2023characteristicsinformed introduced Characteristic-Informed Neural Networks (CINNs) for transport equations, where PDE information is integrated into the neural network architecture. A similar approach, albeit for Dirichlet boundary conditions, was proposed by Lagaris et al. Lagaris_1998 . Our method builds upon these architectural modifications, with a focus on enhancing stability and ensuring proper propagation of information, especially for stiff time-dependent PDEs. We carefully design our architecture to address the challenges of information propagation and loss balancing, aiming to improve the robustness and efficiency of PINN training.
2 Methodology
We consider the following setup for a family of time dependent PDEs. Let $u$ be an unknown function defined on $[0,T]\times\Omega$ with $\Omega\subset\mathcal{R}^{d}$. Here the physical domain $\Omega$ comes with a Lipschitz boundary $\partial\Omega$ and $\bar{\Omega}=\Omega\cup\partial\Omega$. Let $\mathcal{P}$ be an operator which can include differentiation and integration. For example, for the $1D$ heat equation $u_{t}-\lambda u_{xx}=0$, $\mathcal{P}=\lambda\frac{\partial^{2}}{\partial x^{2}}$. Notice that $\mathcal{P}$ can be nonlinear. Let $\mathcal{B}$ be another operator on the boundary, that is, for Dirichlet boundary condition $u(t,x)=g(t,x)$ for $(t,x)\in[0,T]\times\partial\Omega$, $\mathcal{B}=\mathcal{I}$ (the identity). Then we say that $u$ is a solution of an PDE, if $u$ satisfies the following
$\begin{cases}u_{t}-\mathcal{P}(u)(t,\mathbf{x})&=f(t,\mathbf{x}),\quad(t,\mathbf{x})\in(0,T]\times\Omega,\\ u(0,\mathbf{x})&=u_{0}(\mathbf{x}),\quad\mathbf{x}\in\bar{\Omega},\\ \mathcal{B}(u)(t,\mathbf{x})&=g(t,\mathbf{x}),\quad(t,\mathbf{x})\in[0,T]\times\partial\Omega,\\ \end{cases}$ (1)
Here the functions $u_{0}$, $f$, and $g$ are user-inputs and are assumed to satisfy desired regularity so that the existence and uniqueness of solutions for such a PDE is guaranteed. We also assume that the compactibility condition, that is, $g(0,\mathbf{x})=u_{0}(\mathbf{x})$ for $\mathbf{x}\in\partial\Omega$, is satisfied to prevent any ill-conditioning of the PDE problem.
Types of Boundary Conditions: one can consider many kinds of boundary conditions as follows
$\mathcal{B}(u)(t,\mathbf{x})$ $=u(t,\mathbf{x}),\quad\text{Dirichlet type},$
$\mathcal{B}(u)(t,\mathbf{x})$ $=\frac{\partial u}{\partial\mathbf{n}}(t,\mathbf{x}),\quad\text{Neumann type, $\mathbf{n}$ is the outward normal vector to $\partial\Omega$},$
$\mathcal{B}(u)(t,\mathbf{x};a,b)$ $=au(t,\mathbf{x})+b\frac{\partial u}{\partial\mathbf{n}}(t,\mathbf{x}),\quad\text{Robin type}.$
The periodic boundary condition gives $d$ different equations, and they are given as follows,
$\mathcal{B}(u)(t,\mathbf{x})=u(t,\mathbf{x})-u(t,\mathbf{x}+P\mathbf{e}_{i})=0,\quad i=1,\cdots,d.$ (2)
Here $P$ is the period and $\mathbf{e}_{i}$ is the $i^{th}$ standard basis vector in $\mathcal{R}^{d}$, that is,
$\mathbf{e}_{i}=\underbrace{\begin{bmatrix}0&\cdots&0&1&0&\cdots&0\end{bmatrix}^{\top}}_{\text{$1$ is the in the $i^{th}$ position.}}.$
One can even consider a mixture of two or all of the above BC conditions. For example, we can have $\Omega=\Gamma_{1}\cup\cdots\cup\Gamma_{K}$ with each $\Gamma_{i}$ defines a different kind of BC. In this paper, we will mainly focus on the periodic type of boundary conditions, however other kinds of BCs can be
Classical numerical techniques such as the Finite Difference Method (which approximates derivatives on a point mesh), Finite Element Method (utilizing basis approximation on a triangular mesh with weak formulation), and Spectral Method (employing Fourier basis with FFT on a point mesh) are widely utilized in scientific computations. However, recent advancements in scientific machine learning integrate the principles of physics, particularly the understanding of PDEs, into the training process of machine learning models. This integration enables the development of PINNs to solve for $u$ as follows: one tries to find an approximate solution from a set of deep neural networks (of the same depth, same number of neurons on each hidden layer, and same activation functions on each layer) $\mathcal{H}_{\text{NN}}$, which is a minimizer of the following loss functional
$\text{Loss}(u_{nn})=\text{Data Loss}+\lambda*\text{Physical Loss}$
where $\lambda$ is a regularization parameter, and the two losses are as follows
Data $\text{Data Loss}=\frac{1}{N_{IC}}\sum_{i=1}^{N_{IC}}|u_{nn}(0,\mathbf{x}_{i}^{IC})-u_{0}(\mathbf{x}_{i}^{IC})|^{2}$
$\quad+\sum_{i=1}^{d}\frac{1}{N_{BC}}\sum_{j=1}^{N_{BC}}|u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC})-u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC}+P\mathbf{e}_{i})|^{2}$
Physical $\text{Physical Loss}=\frac{1}{N_{CL}}\sum_{i=1}^{N_{CL}}|(\frac{\partial u}{\partial t}-\mathcal{P}(u_{nn})-f)(t_{i}^{CL},\mathbf{x}_{i}^{CL})|^{2}$
for $u_{nn}\in\mathcal{H}_{\text{NN}}$. Here $\{(t^{CL},\mathbf{x}^{CL})_{i}\}_{i=1}^{N_{CL}}\in(0,T]\times\Omega$ are called collocation points, $\{(0,\mathbf{x}^{IC})_{i}\}_{i=1}^{N_{IC}}\in\{0\}\times\bar{\Omega}$ are the initial condition points, and $\{(t^{BC},\mathbf{x}^{BC})_{i}\}_{i=1}^{N_{BC}}\in[0,T]\times\partial\Omega$ are the boundary condition points. The minimizer, denoted as $u_{NN}$, will an be approximate solution to (1).
2.1 Hard Constrains
The main motivation for us to consider hard constrains transformation for PINNs is the possible ill-conditioning of the PDE solution map’s dependence on IC and BC. Traditional numerical methods do not need to encounter this kind of difficulty, as the IC data is exactly satisfied in their solutions. PINNs, on the other hand, use $L^{2}$-loss to fit the IC/BC and PDE residual data, which causes the solutions to be highly sensitive to the tightness of fit for the IC in the training (as pointed out in mcclenny2020self ). By transforming the need of fitting IC/BC conditions out of PINNs, it reduces the spectral bias in the training. Therefore, We consider the following transformation
$\tilde{u}(t,\mathbf{x})=\psi(t,\mathbf{x})+\phi(t,\mathbf{x})u_{nn}(t,\mathbf{x})$
Here the functions $\psi$ and $\phi$ are smooth functions with the following property
$\psi(0,\mathbf{x})=u_{0}(\mathbf{x}),\quad\phi(0,\mathbf{x})=0,\quad\text{for $\mathbf{x}\in\bar{\Omega}$},$
where in the case of periodic BC, we require $\psi$ and $\phi$ to be both $P$-periodic; for other BC types, we instead require $\phi$ and $\psi$ to satisfy
$\psi(t,\mathbf{x})=g(t,\mathbf{x}),\quad\phi(t,\mathbf{x})=0,\quad\text{for $(t,\mathbf{x})\in[0,T]\times\partial\Omega$}.$
When such transformation is used, then the training for $u_{nn}$ is changed, due to the fact that IC/BC can be built into the PDE loss, i.e.
$\text{Loss}(u_{nn})=\frac{1}{N_{CL}}\sum_{i=1}^{N_{CL}}|(\tilde{u}_{t}-\mathcal{P}(\tilde{u}))(t_{i}^{CL},\mathbf{x}_{i}^{CL})|^{2},\quad\text{for $\tilde{u}=\psi+\phi u_{nn}$}.$
We have reduced the types of losses in the training; however the information about IC and BC has then moved into the PDE residual loss through $\psi$ and $\phi$. The difficulty in training the new PDE residual loss has been increased. Hence, other training enhancement techniques can be used to improve the new PDE residual training.
Periodic Neural Networks: even though we require $\psi$ and $\phi$ to be periodic, we still need to build the information of periodicity into $u_{nn}$, hence we consider the following composition
$u_{nn}(t,x)=f_{nn}(v(t,x)),\quad\text{$f_{nn}$ is a neural network, and $x\in\mathcal{R}$},$
and
$v(t,x)=\begin{bmatrix}t&1&\cos(\frac{2\pi}{P}x)&\sin(\frac{2\pi}{P}x)&\cos(\frac{2\pi}{P}2x)&\cdots&\cos(\frac{2\pi}{P}mx)&\sin(\frac{2\pi}{P}mx)\end{bmatrix}^{\top},$
where $m>0$ is a positive integer (another hyper-parameter). With this composition, $u_{nn}$ will automatically become $P$-periodic dong2021method . For higher dimensional physical $\mathbf{x}$, similar transformation can be used.
2.2 Mini-batching
We have eliminated the IC/BC losses, but the transformed PDE loss might create additional difficulties in training due to the introduction of IC/BC functions into the PDE. We adopt a machine learning technique, called mini-batching, commonly used to reduce training complexity by evaluating the loss on randomly selected subsets of collocation data. In PINNs, this reduces memory costs, especially when handling large numbers of collocation points. Sampling mini-batches from collocation points also helps stabilize and accelerate optimization. So we consider mini-batching in the training for decreasing the PDE residual loss. As pointed out in wight2020solving , the mini-batching technique is similar to time-marching resampling. But with mini-batching, one does not employ an explicit time grid.
2.3 Self-Adaptive Physics-Informed Neural Networks
The proper choice of penalty term $\lambda$ can be updated from derived formula as mentioned in wang2020and . However, data-driven discovery of $\lambda$ is possible. Outline in mcclenny2020self , the loss is further modified as
$\text{Data Loss}(\lambda_{IC,i}\}_{i=1}^{N_{IC}},\{\lambda_{BC,i}\}_{i=1}^{N_{BC}})=\frac{1}{N_{IC}}\sum_{i=1}^{N_{IC}}m(\lambda_{IC,i})|u_{nn}(0,\mathbf{x}_{i}^{IC})-u_{0}(\mathbf{x}_{i}^{IC})|^{2}$
$\quad+\sum_{i=1}^{d}\frac{1}{N_{BC}}\sum_{j=1}^{N_{BC}}m(\lambda_{BC,i})|u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC})-u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC}+P\mathbf{e}_{i})|^{2}$
Physical $\text{Physical Loss}(\{\lambda_{PDE,i}\}_{i=1}^{N_{PDE}})=\frac{1}{N_{CL}}\sum_{i=1}^{N_{CL}}m(\lambda_{PDE,i})|(\frac{\partial u}{\partial t}-\mathcal{P}(u_{nn})-f)(t_{i}^{CL},\mathbf{x}_{i}^{CL})|^{2}$
The set of loss weights, $\{\lambda_{IC,i}\}_{i=1}^{N_{IC}}\cup\{\lambda_{BC,i}\}_{i=1}^{N_{BC}}\cup\{\lambda_{PDE,i}\}_{i=1}^{N_{PDE}}$, are designed in a way to increase when needed during any training epoch in order to combat the spectral bias. Adaptive methods are essential to ensure that the neural network accurately addresses the challenging spots in the solutions of “stiff” PDEs. In our experiments, we employ self-adaptive PINNs as proposed by mcclenny2020self to train PINNs adaptively. This approach involves fully trainable adaptation weights applied individually to each training point, allowing the neural network to independently identify and focus on the challenging regions of the solution.
2.4 Performance Measures
Let $\{x_{i},t_{i}\}_{i=1}^{N}$ be a collection of $N$ data points where we evaluate both the reference solution $u(x,t)$ and the neural network output $\mathcal{U}(x,t)$. To measure the accuracy of the trained model, we compute the relative $L_{1}$ norm $\mathcal{E}_{1}$ and relative $L_{2}$ norm $\mathcal{E}_{2}$, defined by:
$\mathcal{E}_{1}=\frac{\sum_{i=1}^{N}|\mathcal{U}(x_{i},t_{i})-u(x_{i},t_{i})|}{\sum_{i=1}^{N}|u(x_{i},t_{i})|},\quad\quad\quad\mathcal{E}_{2}=\frac{\Big{(}\sum_{i=1}^{N}|\mathcal{U}(x_{i},t_{i})-u(x_{i},t_{i})|^{2}\Big{)}^{1/2}}{\sqrt{\sum_{i=1}^{N}|u(x_{i},t_{i})|^{2}}}.$
These metrics represent the normalized differences between the model predictions and the ground truth values, providing a quantitative evaluation of model performance.
3 Numerical Experiments
In this section, we examine the proposed PINN scheme on a diverse set of prototypical stiff time-dependent PDEs with periodic boundary conditions, aiming to demonstrate its robustness and flexibility in handling a broad class of challenging equations. These benchmarks exhibit a wide range of physical and mathematical characteristics: from stiff dynamics (e.g., Kuramoto–Sivashinsky), to multi-species reaction–diffusion systems (e.g., Gray–Scott and Belousov–Zhabotinsky), and nonlinear dispersive waves (e.g., nonlinear Schrödinger equation). Such equations often involve strong nonlinearities, stiffness, and parameter sensitivities, posing significant challenges for data-driven methods such as PINNs. Nonetheless, our results illustrate the capacity of the proposed framework to overcome these difficulties and accurately approximate the solutions across these varied regimes. Detailed experimental results are presented in Section 3.1.
To begin, we employ Latin Hypercube Sampling to generate collocation points across the full spatio-temporal domain, including its boundary. To facilitate training and better incorporate the initial condition into the solution structure, we adopt the following transformation:
$\tilde{u}(t,\mathbf{x})=\exp(-Ct)*u_{0}(\mathbf{x})+t*u_{nn}(t,\mathbf{x}),\quad(t,\mathbf{x})\in[0,T]\times\bar{\Omega};\quad\psi=\exp(-Ct)*u_{0}(\mathbf{x}),\,\phi=t.$
where $u_{0}(\mathbf{x})$ denotes the initial condition, and we test two values of $C\in\{0.1,1\}$, to assess how the decay rate affects learned dynamics. The transformed PINN $u_{nn}$ is trained to minimize a loss function that accounts for the PDE residuals, which now includes IC/BC. To rigorously assess the accuracy of our results, we use Chebfun, a spectral-style computing system in MATLAB that enables highly accurate function approximations. This allows us to make precise comparisons between the ground-truth solutions and the PINN-generated solutions.
For each benchmark, we not only present the main numerical results but also carry out detailed ablation studies to evaluate the contribution of each proposed technique. In each ablation experiment, we deactivate one component of the methodology at a time, keeping the remaining elements active under the same hyper-parameter configurations, to measure the relative $L^{2}$ error and relative $L^{1}$ error. The neural network architecture used throughout consists of 7 hidden layers with 32 neurons per layer and tanh activation functions. The initial weights are fixed across all ablation runs for consistency. Model training proceeds in two stages: first using the Adam optimizer kingma2014adam with a learning rate of $10^{-3}$ for $50,000$ steps, followed by fine-tuning via the L-BFGS-B optimizer with a learning rate of $1$ for $5,000$ steps. Each experiment uses $16,384$ interior collocation points, along with $128$ initial and $128$ boundary data points sampled via LHS. For PDE systems (e.g., Gray–Scott), we further apply a mini-batching strategy with $5$ batches and a batch size of $3,280$ to enhance training stability.
Overall, these extensive experiments highlight the effectiveness and adaptability of our PINN scheme in solving a wide variety of stiff time-dependent PDEs. The results provide not only strong empirical validation but also a reliable and reproducible benchmark for future developments in this area. For full experimental details and visualizations, see Section 3.1.
3.1 Allen-Cahn Equation $I$
The Allen-Cahn type PDE is a classical phase-field model that has been widely used to study phase separation phenomena bazant2017thermodynamic . It has numerous practical applications across a range of fields, including materials science allen1979microscopic ; shen2010numerical , biological systems hyman2014liquid ; takatori2015towards , and electrochemical systems horstmann2013precipitation ; tian2015electrochemically . We begin with the one-dimensional Allen-Cahn equation with periodic boundary conditions, a representative example that presents challenges for traditional PINN models, formulated as follows:
$u_{t}-\gamma_{1}u_{xx}+\gamma_{2}u^{3}-\gamma_{2}u=0,\quad(t,x)\in(0,T)\times(a,b),$ (3)
$u(0,x)=u_{0}(x),\quad x\in[a,b],$
$u(t,a)=u(t,b),\quad t\in[0,T],$
$u_{x}(t,a)=u_{x}(t,b),\quad t\in[0,T],$
where $\gamma_{1},\gamma_{2}>0,T>0,a<b$ are prescribed constants. As $\gamma_{2}$ increases, the transition interface of the solutions is sharper, which makes it harder to solve the AC equation numerically. We first demonstrate the effectiveness of our scheme by testing on Allen-Cahn equation with the following initial condition and parameters, $u_{0}(x)=x^{2}\cos(\pi x)$, $T=1$, $a=-1$, $b=1$, $\gamma_{1}=0.001$, and $\gamma_{2}=5$.
Figure 1: Exact solution of the 1D Allen-Cahn Equation (Case I) with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $1.18\times 10^{-3}$ and the relative $L^{1}$ error is $5.34\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Figure 2: Best solutions of the Allen-Cahn Equation (Case I) $\gamma_{2}=5$. The resulting relative $L^{2}$ error is $1.18\times 10^{-3}$ and the relative $L^{1}$ error is $5.34\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
For this example, we first test the baseline PINN approach raissi2019physics . As shown in Table 1, the baseline PINN alone fails to accurately solve the Allen-Cahn equation, motivating the development of a more effective training pipeline. To address this, we employ all proposed techniques (except mini-batching) and conduct an ablation study by disabling individual components one at a time. Additionally, we introduce two different decay functions for the initial condition, $e^{-1t}$ and $e^{-0.1t}$, to investigate their influence on the solution. The results are summarized in Table 1. The full scheme, including self-adaptive loss balancing, achieves the best performance, reaching a relative $L^{2}$ error of $1.18\times 10^{-3}$ and relative $L^{1}$ error of $5.34\times 10^{-4}$. In particular, using the faster-decaying initial condition ($IC1$) leads to better results than the slower-decaying case ($IC2$), suggesting that the initial condition plays a more critical role during the early stages of training but becomes less dominant over time. The ablation results further highlight the importance of each methodological component. Removing any single part generally leads to performance degradation, with the most severe impact observed when the initial condition is omitted, resulting in a relative error close to $0.99$ in both $L^{1}$ and $L^{2}$ norms. Figure 1 illustrates the exact solution, the network prediction, and the absolute error distribution. As seen, the predicted solution matches the ground truth remarkably well, with errors concentrated at negligible levels. Figure 2 presents cross-sectional comparisons at different times, further confirming the excellent agreement between the predicted and true solutions across the entire time horizon. These results provide strong evidence for the effectiveness of the proposed training strategy and underscore the critical role of each pipeline component.
Table 1: Allen-Cahn Equation (Casse $I$ and Case $II$): Relative $L^{2}$ and $L^{1}$ errors for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
Ablation Settings Performance - Case $I$ Performance - Case $II$
PBC IC1 IC2 SA Rel. $L^{2}$ error Rel. $L^{1}$ error Rel. $L^{2}$ error Rel. $L^{1}$ error
$1.18\times 10^{-3}$ $5.34\times 10^{-4}$ $3.53\times 10^{-4}$ $2.14\times 10^{-4}$
$7.64\times 10^{-2}$ $1.66\times 10^{-2}$ $7.04\times 10^{-4}$ $3.26\times 10^{-4}$
$1.24\times 10^{-3}$ $6.10\times 10^{-4}$ $8.32\times 10^{-4}$ $3.72\times 10^{-4}$
$2.86\times 10^{-2}$ $8.20\times 10^{-3}$ $9.80\times 10^{-4}$ $3.91\times 10^{-4}$
$9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$
$9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$
$1.05\times 10^{-2}$ $2.65\times 10^{-3}$ $5.11\times 10^{-2}$ $1.21\times 10^{-2}$
$1.86\times 10^{-3}$ $7.31\times 10^{-4}$ $5.03\times 10^{-2}$ $8.27\times 10^{-3}$
$1.88\times 10^{-2}$ $4.71\times 10^{-3}$ $4.58\times 10^{-2}$ $1.10\times 10^{-2}$
$4.34\times 10^{-3}$ $1.24\times 10^{-3}$ $3.03\times 10^{-2}$ $4.25\times 10^{-3}$
$5.12\times 10^{-1}$ $3.18\times 10^{-1}$ $4.02\times 10^{-1}$ $1.42\times 10^{-1}$
$9.99\times 10^{-1}$ $9.98\times 10^{-1}$ $5.11\times 10^{-1}$ $3.28\times 10^{-1}$
3.2 Allen-Chan, Case $II$
For our second example, we investigate the performance of PINNs in scenarios involving sharper moving interfaces. Specifically, we consider a modified Allen–Cahn equation with more oscillatory initial data:
$u_{t}-\gamma_{1}u_{xx}+\gamma_{2}u^{3}-\gamma_{2}u=0,\quad(t,x)\in(0,1)\times(-1,1),$ (4)
$u(0,x)=x^{2}\sin(2\pi x),\quad x\in[-1,1],$
$u(t,-1)=u(t,1),\quad t\in[0,1],$
$u_{x}(t,-1)=u_{x}(t,1),\quad t\in[0,1],$
where $\gamma_{1}=0.001,\gamma_{2}=4$.
We first train the baseline PINN model, which produces a large relative $L^{2}$ error of $0.511$ and relative $L^{1}$ error of $0.328$, demonstrating its inadequacy in resolving the sharp interface dynamics (Table 1). However, once the initial condition enforcement and periodic boundary neural networks are incorporated, the performance improves drastically, achieving a relative $L^{2}$ error of $8.32\times 10^{-4}$ and relative $L^{1}$ error of $3.72\times 10^{-4}$. Further gains are realized when using the self-adaptive loss weighting strategy, although the improvement is marginal compared to the dominant contribution from enforcing the initial and boundary conditions. Interestingly, unlike the previous case, there is minimal performance difference between the schemes using $IC1$ and $IC2$, suggesting that the impact of initial condition decay becomes less significant when strong oscillations are already present at the beginning.
Figure 3: Exact solution of the Allen-Cahn Equation (Case II) with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $3.53\times 10^{-4}$ and the relative $L^{1}$ error is $2.14\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Figure 4: Best solutions of the Allen-Cahn Equation (Case II) $\gamma_{2}=4$. The resulting relative $L^{2}$ error is $3.53\times 10^{-4}$ and the relative $L^{1}$ error is $2.14\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
3.3 Cahn-Hilliard Equation
Figure 5: Exact solution of the 1D Cahn-Hilliard Equation with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $9.75\times 10^{-4}$ and the relative $L^{1}$ error is $5.65\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Figure 6: Best solutions of the 1D Cahn-Hilliard Equation. The resulting relative $L^{2}$ error is $3.53\times 10^{-4}$ and the relative $L^{1}$ error is $2.14\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
The Cahn-Hilliard equation, a variant of the Allen-Cahn equation, also models phase separation. This process describes how the components of a binary fluid spontaneously separate into distinct domains, each composed of a single component cahn1958free ; kim2016basic ; lee2014physical ; miranville2017cahn . The equation is expressed as:
$u_{t}=\epsilon_{1}(-u_{xx}-\epsilon_{2}u_{xxxx}+(u^{3})_{xx}),\quad(t,x)\in(0,T]\times(-L,L),$ (5)
$u(0,x)=u_{0}(x),\quad x\in[-L,L],$
$u(t,-L)=u(t,L),\quad t\in[0,T],$
where $\epsilon_{1}=10^{-2},\epsilon_{2}=10^{-4},T=1,L=1$. We specify the initial condition as $u_{0}(x)=-\cos(2\pi x)$. This PDE involves higher-order derivatives, making it more challenging to solve compared to the Allen-Cahn equation.
For the ablation study in Section 2, the neural network is trained using the same hyperparameters outlined in Section 3. In this experiment, the full scheme yields the best results. Additionally, methods using initial condition IC2 outperform those with IC1. Except for the full schemes and the experiment with periodic neural networks for boundary conditions and initial condition IC2, none of the other approaches produce accurate solutions. The errors for all experiments are summarized in Table 2. The best results are illustrated in Figures 5 and 6.
Table 2: Cahn-Hilliard Equation: Relative $L^{2}$ and $L^{1}$ errors for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
Ablation Settings Performance
PBC IC1 (C=1) IC2 (C=0.1) SA Rel. $L^{2}$ error Rel. $L^{1}$ error
$1.06\times 10^{-3}$ $5.61\times 10^{-4}$
$9.75\times 10^{-4}$ $5.65\times 10^{-4}$
$2.50\times 10^{-1}$ $5.87\times 10^{-2}$
$3.87\times 10^{-3}$ $1.91\times 10^{-3}$
$9.99\times 10^{-1}$ $9.99\times 10^{-1}$
$9.99\times 10^{-1}$ $9.99\times 10^{-1}$
$3.48\times 10^{-1}$ $2.94\times 10^{-1}$
$2.37\times 10^{-1}$ $1.09\times 10^{-1}$
$4.07\times 10^{-1}$ $3.45\times 10^{-1}$
$1.55\times 10^{-2}$ $8.20\times 10^{-3}$
$9.99\times 10^{-1}$ $9.99\times 10^{-1}$
$9.99\times 10^{-1}$ $9.98\times 10^{-1}$
3.4 Kuramoto-Sivashinsky Equation
Kuramoto–Sivashinsky equation is a fourth-order nonlinear partial differential equation, known for its chaotic behavior kudryashov1990exact ; michelson1986steady . It has served as a pivotal model in the study of complex dynamical phenomena observed in various physical systems. Consequently, the Kuramoto–Sivashinsky equation provides an excellent case study to demonstrate how PINNs can effectively simulate chaotic dynamics. The equation is given by:
$u_{t}=-u_{xx}-u_{xxxx}-uu_{x},\quad(t,x)\in(0,T]\times(a,b),$ (6)
$u(0,x)=u_{0}(x),\quad x\in[a,b],$
$u(t,a)=u(t,b),\quad t\in[0,T],$
where $T=20,a=0,b=32\pi$, and $u_{0}(x)=\cos\frac{x}{16}(1+\sin\frac{x-1}{16})$. The parameters are the same as those mentioned in Section 3. The best results are shown in Figure 7 and 8:
Figure 7: Exact solution of the 1D Kuramoto-Sivashinsky Equation with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $1.46\times 10^{-3}$ and the relative $L^{1}$ error is $1.02\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 8: Best solutions of the 1D Kuramoto-Sivashinsky Equation. The resulting relative $L^{2}$ error is $1.46\times 10^{-3}$ and the relative $L^{1}$ error is $1.02\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Table 3: Kuramoto-Sivashinsky Equation: Relative $L^{2}$ and $L^{1}$ errors for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
Ablation Settings Performance
PBC IC1 (C=1) IC2 (C=0.1) SA Rel. $L^{2}$ error Rel. $L^{1}$ error
$4.04\times 10^{-2}$ $1.46\times 10^{-2}$
$1.46\times 10^{-3}$ $1.02\times 10^{-3}$
$1.68\times 10^{-2}$ $8.94\times 10^{-3}$
$1.30\times 10^{-2}$ $7.79\times 10^{-3}$
$9.84\times 10^{-3}$ $6.36\times 10^{-3}$
$2.50\times 10^{-3}$ $1.81\times 10^{-3}$
$2.13\times 10^{-1}$ $9.74\times 10^{-2}$
$2.00\times 10^{-1}$ $1.04\times 10^{-1}$
$1.54\times 10^{-2}$ $1.21\times 10^{-2}$
$9.60\times 10^{-3}$ $8.75\times 10^{-3}$
$2.11\times 10^{-1}$ $1.06\times 10^{-1}$
$3.01\times 10^{-1}$ $1.97\times 10^{-1}$
As in Table 3, solving this PDE system using a baseline PINN alone proves to be exceptionally challenging. The incorporation of periodic neural networks for boundary conditions, paired with initial condition IC2, consistently yields superior results compared to IC1. This suggests that, in contrast to the Allen-Cahn equation discussed in Section 3.1, the choice of initial condition plays a more significant role in the training process for this more complex system. This observation underscores the importance of carefully selecting both the boundary conditions and initial conditions when training neural networks for chaotic or higher-order PDE systems, where these factors contribute notably to the accuracy and stability of the solution.
3.5 Gray-Scott Equation
Reaction and diffusion of chemical species can produce a variety of patterns de1991turing ; nishiura1999skeleton , reminiscent of those often seen in nature. The Gray-Scott type system is one of classical mathematical models for chemical reactions gray1983autocatalytic ; gray1984autocatalytic ; liang2022reversible . The general irreversible Gray-Scott equations describe such reactions:
$U+2V$ $\longrightarrow 3V,$ (7)
$V$ $\longrightarrow P.$
This system is defined by two equations that describe the dynamics of two reacting substances:
$u_{t}=\epsilon_{1}u_{xx}+b(1-u)-uv^{2},\quad(t,x)\in(0,T]\times(-L,L),$ (8)
$v_{t}=\epsilon_{2}v_{xx}-(b+k)v+uv^{2},$
$u(0,x)=u_{0}(x),\quad v(0,x)=v_{0}(x),\quad\forall x\in[-L,L],$
$u(t,-L)=u(t,L),\quad v(t,-L)=v(t,L),\quad\forall t\in[0,T],$
where $T=20,L=50$, $\epsilon_{1}=1,\epsilon_{2}=0.01$ are diffusion rates, $b=0.02$ is the "feeding rate" that adds $U$, $k=0.0562$ is the "killing rate" that removes $V$. We set our initial conditions as:
$u_{0}(x)=1-\frac{\sin(\pi(x-50)/100)^{4}}{2},\quad v_{0}(x)=\frac{\sin(\pi(x-50)/100)^{4}}{4}.$
In this ablation study, we not only evaluate the performance of self-adaptive neural networks but also test our scheme with mini-batching on the Gray-Scott equations, which model the densities of two interacting species. For mini-batching, we use 5 batches, each containing 3,280 samples. As shown in Table 4, the results confirm that each proposed component improves the overall model performance. Omitting any one of these components leads to higher error rates. Notably, the errors for the $u-$equation are typically smaller than those for the $v-$equation, which aligns with our expectations. This is because the $v-$equation is sharper and more challenging to train compared to the $u-$equation. Additionally, experiments using initial condition IC2 consistently outperform those using IC1. Specifically, the full scheme with IC2 achieves the best results, with a relative $L^{2}$ error of $2.05\times 10^{-4}$ for $u-$equation and $8.46\times 10^{-4}$ for $v-$equation. These results suggest that IC2 is more effective for larger time-scale PDEs, as IC1 causes the initial condition to decay too quickly, reducing its influence during training. The predicted solutions from our best model are visualized in Figures 9 and 10, respectively.
Figure 9: Best solutions of the 1D Gray-Scott Equation (u species). The resulting relative $L^{2}$ error is $2.05\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Figure 10: Best solutions of the 1D Gray-Scott Equation (v species). The resulting relative $L^{2}$ error is $2.05\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Figure 11 and Figure 12 show the comparison of the best prediction against the reference solution. Note that the majority of this error is attributed to the sharpness of the PDEs.
Figure 11: Exact solution of the 1D Gray-Scott Eqaution (v specie) with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $8.46\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Figure 12: Exact solution of the 1D Gray-Scott Eqaution (v specie) with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $8.46\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Table 4: Gray-Scott Equation: Relative $L^{2}$ errors of $u$ and $v$ equations for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
Ablation Settings Performance
PBC IC1 IC2 SA mini-batch Rel. $L^{2}$ error (u - eq.) Rel. $L^{2}$ error (v - eq.)
$8.98\times 10^{-4}$ $7.92\times 10^{-3}$
$2.05\times 10^{-4}$ $8.46\times 10^{-4}$
$4.93\times 10^{-3}$ $3.03\times 10^{-2}$
$2.86\times 10^{-3}$ $8.20\times 10^{-3}$
$9.01\times 10^{-3}$ $4.27\times 10^{-3}$
$7.59\times 10^{-3}$ $3.98\times 10^{-3}$
$5.11\times 10^{-3}$ $2.58\times 10^{-2}$
$1.23\times 10^{-3}$ $9.74\times 10^{-3}$
$8.09\times 10^{-1}$ $9.99\times 10^{-1}$
$2.06\times 10^{-1}$ $7.43\times 10^{-1}$
$5.37\times 10^{-1}$ $2.58\times 10^{-1}$
$1.15\times 10^{-1}$ $2.62\times 10^{-1}$
$9.53\times 10^{-2}$ $8.65\times 10^{-2}$
$9.21\times 10^{-2}$ $2.98\times 10^{-1}$
3.6 Belousov-Zhabotinsky Equation
A Belousov–Zhabotinsky reaction is one of a class of reactions that serve as a classical example of non-equilibrium thermodynamics. It typically involves the oxidation of organic compounds by bromine in an acidic medium zhabotinsky2007belousov ; cassani2021belousov . The system of PDEs takes the following form:
$u_{t}=\epsilon_{1}u_{xx}+u+v-uv-u^{2},$ (9)
$v_{t}=\epsilon_{2}v_{xx}+w-v-uv,$
$w_{t}=\epsilon_{1}w_{xx}+u-w,\quad(t,x)\in(0,T]\times[-L,L]$
$u(0,x)=u_{0}(x),\quad v(0,x)=v_{0}(x),\quad w(0,x)=w_{0}(x),\quad\forall x\in[-L,L]$
$u(t,-L)=u(t,L),\quad v(t,-L)=v(t,L),\quad w(t,-L)=w(t,L),\quad\forall t\in[0,T],$
where $T=2,L=1$, $\epsilon_{1}=10^{-5},\epsilon_{2}=2\times 10^{-5}$ are diffusion rates. We let the initial conditions be:
$u_{0}(x)=\exp(-100(x+0.5)^{2}),\quad v_{0}(x)=\exp(-100x^{2}),\quad w_{0}(x)=\exp(-100(x-0.5)^{2}).$
The system of PDEs involving three species provides a valuable opportunity to evaluate our proposed schemes using the techniques described in Section 2 for simulating multispecies dynamics. The results, summarized in Table 5, show that the full scheme consistently achieves the best overall performance, in line with previous experiments. Interestingly, the errors observed for methods using IC1 and those using IC2 are not significantly different. Both IC1 and IC2, when used independently, are capable of producing satisfactory results. However, when periodic neural networks are used to enforce boundary conditions, the performance declines and becomes less stable, even though self-adaptive training and mini-batching provide some improvements. This observation suggests that incorporating boundary conditions introduces additional complexity into the training process. Further investigation into their effects would be a promising direction for future research.
Figure 13: Best solutions of the 1D Belousov-Zhabotinsky equation (u specie). The resulting relative $L^{2}$ error is $3.49\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 14: Best solutions of the 1D Belousov-Zhabotinsky equation (v specie). The resulting relative $L^{2}$ error is $2.45\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 15: Best solutions of the 1D Belousov-Zhabotinsky equation (w specie). The resulting relative $L^{2}$ error is $2.50\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
The comparisons between the exact solutions of $(u,v,w)$ equations with corresponding best neural network predictions on the whole $(x,t)$ domain are shown in Figures 16, 17 and 18. Although there is a slight misalignment between each two of them, the model predictions closely match the corresponding numerical estimations.
Figure 16: Exact solution of the 1D Belousov-Zhabotinsky equation (u specie) with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $3.49\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 17: Exact solution of the 1D Belousov-Zhabotinsky equation (v specie) Equation with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $2.45\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 18: Exact solution of the 1D Belousov-Zhabotinsky equation (w specie) Equation with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $2.50\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Table 5: Belousov-Zhabotinsky Equation: Relative $L^{2}$ errors of $u$, $v$ and $w$ equations for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
Ablation Settings Performance: relative $L^{2}$ error
PBC IC1 IC2 SA mini-batch u-equation v-equation w-equation
$3.49\times 10^{-3}$ $2.45\times 10^{-3}$ $2.50\times 10^{-3}$
$6.06\times 10^{-3}$ $4.47\times 10^{-3}$ $4.30\times 10^{-3}$
$7.08\times 10^{-3}$ $4.77\times 10^{-3}$ $4.24\times 10^{-3}$
$5.03\times 10^{-3}$ $3.10\times 10^{-3}$ $3.71\times 10^{-3}$
$9.17\times 10^{-3}$ $9.29\times 10^{-3}$ $8.89\times 10^{-3}$
$1.05\times 10^{-1}$ $1.78\times 10^{-2}$ $1.29\times 10^{-2}$
$2.22\times 10^{-2}$ $1.33\times 10^{-2}$ $1.52\times 10^{-2}$
$1.54\times 10^{-2}$ $9.17\times 10^{-3}$ $1.07\times 10^{-2}$
$9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$
$7.99\times 10^{-3}$ $5.05\times 10^{-3}$ $5.16\times 10^{-3}$
$9.39\times 10^{-3}$ $9.41\times 10^{-3}$ $9.85\times 10^{-3}$
$3.98\times 10^{-2}$ $2.15\times 10^{-2}$ $2.79\times 10^{-2}$
$4.48\times 10^{-2}$ $2.71\times 10^{-2}$ $3.22\times 10^{-2}$
$7.87\times 10^{-2}$ $3.95\times 10^{-2}$ $5.48\times 10^{-2}$
3.7 Nonlinear Schröedinger Equation
In theoretical physics, nonlinear Schröedinger equation is a nonlinear PDE, applicable to classical and quantum mechanics kato1987nonlinear ; kevrekidis2001discrete ; fibich2015nonlinear . The dimensionless equation of the classical field is
$u_{t}=iu_{xx}+i|u|^{2}u,\quad(t,x)\in[0,2]\times[-\pi,\pi],$ (10)
$u(0,x)=u_{0}(x),\quad x\in[-\pi,\pi],$
$u(t,-\pi)=u(t,\pi),\quad t\in[0,2],$
where we let
$u_{0}(x)=\frac{2}{2-\sqrt{2}\cos(x)}-1.$
Using the same parameter settings and training steps with prior experiments, we obtain the following results:
Figure 19: Exact solution of the real part of 1D Nonlinear Schroedinger Equation with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $1.80\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 20: Best real-part solutions of the 1D Nonlinear Schroedinger Equation. The resulting relative $L^{2}$ error is $1.80\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 21: Exact solution of the imaginary part of 1D Nonlinear Schroedinger Equation with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $2.58\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 22: Best imaginary-part solutions of the 1D Nonlinear Schroedinger Equation. The resulting relative $L^{2}$ error is $2.58\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Table 6 presents the results of all experiments. The full scheme achieves the best performance, with a relative $L^{2}$ error of real part of $1.80\times 10^{-4}$ and the relative $L^{2}$ error of imaginary part of $2.58\times 10^{-4}$. There is no significant difference in the errors between methods using IC1 and those using IC2. Moreover, without the use of self-adaptive neural networks or mini-batching, accurately obtaining the solution becomes challenging, highlighting the importance of both methods in ensuring more stable training.
Table 6: Nonlinear Schroedinger Equation: Relative $L^{2}$ errors of real and imaginary equations for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
Ablation Settings Performance: relative $L^{2}$ error
PBC IC1 IC2 SA mini-batch Real part Imaginary part
$1.80\times 10^{-3}$ $2.58\times 10^{-3}$
$1.84\times 10^{-3}$ $3.27\times 10^{-3}$
$7.34\times 10^{-2}$ $1.23\times 10^{-1}$
$8.32\times 10^{-2}$ $1.39\times 10^{-1}$
$1.10\times 10^{-2}$ $2.10\times 10^{-2}$
$1.70\times 10^{-2}$ $3.27\times 10^{-2}$
$1.00\times 10^{-1}$ $1.83\times 10^{-1}$
$9.56\times 10^{-2}$ $1.56\times 10^{-1}$
$8.09\times 10^{-1}$ $9.99\times 10^{-1}$
$9.99\times 10^{-1}$ $8.78\times 10^{-1}$
$5.04\times 10^{-2}$ $1.21\times 10^{-1}$
$5.44\times 10^{-2}$ $7.26\times 10^{-2}$
$6.60\times 10^{-2}$ $1.05\times 10^{-1}$
$3.63\times 10^{-1}$ $5.06\times 10^{-1}$
4 Conclusion
We provide an ablation study of hard-constraints and self-adaptive loss weights to study the solution dependency of time-dependent PDEs on its initial conditions. The transformation provided by hard-constraints significantly alleviates the training challenges typically associated with stiff time-dependent PDEs subject. Our study demonstrates effectiveness across seven distinct types of such PDEs, encompassing scalar-valued outputs, vector-valued outputs, and even complex-valued solutions. This approach not only enhances training efficiency but also improves the accuracy and stability of the learned solutions, particularly in the presence of stiffness, where conventional PINNs often struggle to converge.
Although we provide an ablation study of hard-constraints and self-adaptive loss weights, our list of possible training remedies for PINNs is not exhaustive. Moreover, our tests should also include elliptical PDEs with high oscillatory coefficients. Future work includes extending the hard-cosntraint framework to two- and three-dimensional PDEs, which pose additional challenges in both architecture design and computational cost. Ongoing research also focuses on identifying optimal transformations $\psi$ and $\phi$, which govern the embedding of IC and BC into the network structure, with the goal of further enhancing performance and generalizability.
Acknowledgements
BH and MZ thank Prof. Yannis Kevrekids (JHU) for the discussion of well-conditioning of PDE solution maps, and Prof. George Karniadakis (Brown) for his suggestion on training enhancement methods for PINNs. BH developed and implemented the algorithm, ran the simulation and analyzed the data. UB and MZ designed the reserach. All authors wrote the paper. MZ is supported by NSF-AoF grant #$\#2225507$ and Ralph E. Powe Junior Faculty Enhancement Award from ORAU for FY24242424.
Supplementary Material
The source code, initial training data, initial weights, Chebfun code, and ground truth solutions used for benchmarking will provided in https://github.com/baolihao/Training-PINNs-with-Hard-Constraints-and-Adaptive-Weights. The list of hyperparameters and computational resources used to produce the results will be discussed later in this appendix.